### v1

@@ Acknowledgments
Appendix A: Errors in observations
As explained in Section 2, there is clearly an effect where resampling of original observation data to, e.g., hourly temporal resolution will introduce small errors in the tidal analysis, and hence result in outliers when subtracting the computed tidal signal from the observed TWL to obtain the storm surge. An example of this feature is shown in Figure 19, where we show a comparison of the the storm surge signal for the (randomly selected) station Aberdeen (UK) extracted from the observed TWL from the national service British Oceanographic Data Centre (BODC) and the Pan-European CMEMS.
@@
Figure 19: Time interpolation of the observed TWL can create a noisy storm surge signal when subtracting the tides resulting from a harmonic analysis from the TWL. To show this feature we have selected the station of Aberdeen in the United Kingdom. In the GESLA3 dataset this station is represented by a record originating from the British Oceanographic Data Centre (BODC) with a time resolution of combined hourly (pre 1993) and 15 minutes resolution, and a record from CMEMS with hourly resolution. When inspecting the two, the latter seems to utilize some kind of time interpolation based on the first. The resulting storm surge signal is displayed in the time series in panels (a) and (b) based on the BODC and CMEMS data, respectively. An example of how this influence the validation is shown in panels (c) and (d), where the heatmap scatter plots (on log-scale) show the comparison between the model (Diagnostic IB) and observations based on the BODC and CMEMS data, respectively.
Appendix B: Formulas
The formulas used to calculate the error statistics are given in Equations 1, 2, 3 and 4.
@@
$\mbox{MAE}/\sigma=\frac{\sum_{i=1}^{n}\mid\frac{\mbox{obs}_{i}-\mbox{mod}_{i}}{\sigma}\mid}{n},$ (3)
$\mbox{RMSE}/\sigma=\sqrt{\frac{\sum_{i=1}^{n}(\frac{\mbox{obs}_{i}-\mbox{mod}_{i}}{\sigma})^{2}}{n}}.$ (4)
Here, the standard deviation of the observations is given as
$\sigma=\frac{1}{n-1}\sqrt{\sum_{i=1}^{n}(\mbox{obs}_{i}-\overline{\mathrm{obs}}_{i}})^{2}.$ (5)

### v2

@@ Acknowledgments
Appendix A Errors in observations
As explained in Section 2, there is clearly an effect where resampling of original observation data to, e.g., hourly temporal resolution will introduce small errors in the tidal analysis, and hence result in outliers when subtracting the computed tidal signal from the observed TWL to obtain the residual water level. An example of this feature is shown in Figure 19, where we show a comparison of the the residual water level signal for the (randomly selected) station Aberdeen (UK) extracted from the observed TWL from the national service British Oceanographic Data Centre (BODC) and the Pan-European CMEMS.
@@ Appendix A Errors in observations
Figure 19: Time interpolation of the observed TWL can create a noisy residual water level signal when subtracting the tides resulting from a harmonic analysis from the TWL. To show this feature we have selected the station of Aberdeen in the United Kingdom. In the GESLA3 dataset this station is represented by a record originating from the British Oceanographic Data Centre (BODC) with a time resolution of combined hourly (pre 1993) and 15 minutes resolution, and a record from CMEMS with hourly resolution. When inspecting the two, the latter seems to utilize some kind of time interpolation based on the first. The resulting residual water level signal is displayed in the time series in panels (a) and (b) based on the BODC and CMEMS data, respectively. An example of how this influence the validation is shown in panels (c) and (d), where the heatmap scatter plots (on log-scale) show the comparison between the model (Diagnostic IB) and observations based on the BODC and CMEMS data, respectively.
Appendix B Formulas
The quantities used to assess the error statistics are the mean absolute error (MAE),
@@ Appendix B Formulas
as well as their scaled counterparts, $\mbox{MAE}/s$ and $\mbox{RMSE}/s$. Here, the sample standard deviation is computed as
$s=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(\mbox{obs}_{i}-\overline{\mathrm{obs}})^{2}}$ (3)
where sample means are indicated by overbars and computed as
$\overline{\mathrm{obs}}=\frac{1}{n}\sum_{i=1}^{n}\mbox{obs}_{i}.$ (4)